Click Submit to complete this assessment. Question 2 Using the tables for water determine the specified property data at the indicated state. For H20 at p 5 MPa and T 165 °C, find U in kJ/kg. Note: Round up the answer to 2 decimal places. Click Submit to complete this assessment. 5
Answer #1

Pressure = P = 5 Mpa

temperature = T = 165 C

Internal energy = U = ?

Internal energy = U = C*dT

Internal energy = U = 4.2 * 165 = 693 KJ/Kg-K